# Low-Frequency Black-Box Backdoor Attack via Evolutionary Algorithm

### 23 Feb 2024 v1

@@ 5 Conclusion
In this work, we propose a robust black-box backdoor attack by inserting imperceptible perturbations in the frequency domain. Our method provides perceptual similarity in the pixel domain and imperceptibility in the frequency domain. We leverage SA to optimize the trigger to satisfy the attack effectiveness and dual-space stealthiness. The empirical experiments demonstrate that LFBA can achieve a practical attack robustness to elude SOTA defenses in both spatial and frequency domains as well as image processing-based defenses.
Discussion and Limitation. In this work, we concentrate on various computer vision tasks, which have been the focus of numerous existing works [36, 40, 10, 11, 49]. In the future, we intend to expand the scope of this work to other SOTA model architectures, e.g., diffusion and transformer models. The trigger search process is executed in a hybrid GPU-CPU environment during trigger evaluation and optimization phases. It deserves further efforts to design a GPU-accelerated SA to minimize data transmission across hardware, thus improving the efficiency of our proposed LFBA. To counter the proposed attacks, we hope to devise more robust frequency defenses that can surpass the current assumption of both spatial and frequency domain-based backdoor attacks. We hope this work can inspire follow-up studies that provide more secure and robust machine learning systems against our LFBA attack through a frequency perspective.
@@ 0.A.1 Details of Computer Vision Tasks
0.A.2 More Implementation Details
The implementation of our attack LFBA is based on PyTorch [37] and executed on a workstation with 16-core AMD Ryzen 9 7950X CPU, NVIDIA GeForce RTX 4090 and 64G RAM.
@@ 0.A.2 More Implementation Details
0.A.3 Poison Ratio $\rho$
$\rho$ is the fraction of poisoned samples in the training dataset of the adversary. We test the attack effectiveness under different $\rho$ varies from 0.1% to 10%. Although we increase $\rho$ from a wide range, LFBA does not harm the ASR of the victim models. As stated in Figure 8, this fraction setting cannot degrade the ACC and meanwhile, we would like to explore further to examine the lower bound of the fraction which LFBA’s effectiveness can tolerate. Although $\rho$ is $0.1\%$, LFBA can still provide a high ASR, around 80% for GTSRB. We also find that sensitivities to poison ratio can vary among tasks. In CIFAR-10, LFBA achieves above $86\%$ ASR under $\rho=0.5\%$ while it drops rapidly, around $20\%$ when $\rho$ reduce to $0.1\%$.
@@ 0.A.3 Poison Ratio $\rho$
0.A.4 Explanation of Robustness through Frequency Perspective.
We showcase poisoned images and their frequency disparities (compared to clean images) under these image transformations in Figure 9. We can first observe that the frequency disparities of BadNets remain similar to the original ones after JPEG compression while the Gaussian filter destroys the BadNets patterns on both datasets. This proves the fact, as shown in Table 3, that BadNets is effective against JPEG compression but fails to survive after Gaussian filtering. For FTrojan and LFBA-Full, we cannot see any frequency patterns after these transformations on both datasets. However, the frequency disparities of LFBA-Low can be clearly seen even after such operations which indicates our low-frequency attack is robust against preprocessing-based defenses. The reason is low-frequency components are more resilient to such image transformations compared to mid- and high-frequency ones.

### 06 Mar 2024 v2

@@ 5 Conclusion
We propose a robust black-box backdoor attack by inserting imperceptible perturbations in the frequency domain. Our method provides perceptual similarity in the pixel domain and imperceptibility in the frequency domain. We leverage SA to optimize the trigger to satisfy the attack effectiveness and dual-space stealthiness. The empirical experiments demonstrate that LFBA can achieve a practical attack robustness to evade SOTA defenses in both spatial and frequency domains as well as image preprocessing-based defenses.
Discussion and Limitation. In this work, we concentrate on various computer vision tasks, which have been the focus of numerous existing works [36, 40, 10, 11, 49]. In the future, we intend to expand the scope of this work to other SOTA model architectures, e.g., diffusion and transformer models. The trigger search process is executed in a hybrid GPU-CPU environment during trigger evaluation and optimization phases. It deserves further efforts to design a GPU-accelerated SA to minimize data transmission across hardware, thus improving the efficiency of our proposed LFBA. To counter the proposed attacks, we hope to devise more robust frequency defenses that can surpass the current assumption of both spatial and frequency domain-based backdoor attacks. We also hope this work can inspire follow-up studies that enhance the security and robustness of machine learning systems against LFBA through a frequency perspective.
@@ 0.A.1 Details of Computer Vision Tasks
0.A.2 Poison Ratio $\rho$
$\rho$ is the fraction of poisoned samples in the training dataset of the adversary. We test the attack effectiveness under different $\rho$ varying from 0.1% to 10%. Although we increase $\rho$ from a wide range, LFBA does not harm the ASR of the victim models. As stated in Figure 8, this fraction setting cannot degrade the ACC and meanwhile, we would like to examine the lower bound of the fraction that LFBA’s effectiveness can withstand. Even when $\rho$ is $0.1\%$, LFBA can still provide a high ASR, around 80% for GTSRB. We also find that sensitivities to poison ratio can vary among tasks. In CIFAR-10, LFBA achieves above $86\%$ ASR under $\rho=0.5\%$ while it drops rapidly, around $20\%$, when $\rho$ reduces to $0.1\%$.
@@ 0.A.2 Poison Ratio $\rho$
0.A.3 Explanation of Robustness through Frequency Perspective.
We showcase poisoned images and their frequency disparities (compared to clean images) under the image transformations in Figure 9. We can see that the frequency disparities of BadNets remain similar to the original ones after JPEG compression while the Gaussian filter destroys the BadNets patterns on both datasets. This proves the fact, as shown in Table 3, that BadNets is effective against JPEG compression but fails to survive after Gaussian filtering. For FTrojan and LFBA-Full, we cannot see any frequency patterns after these transformations. However, the frequency disparities of LFBA-Low can be clearly seen even after such operations, indicating our low-frequency attack is robust against preprocessing-based defenses. We note that low-frequency components exhibit greater resilience to image transformations than mid- and high-frequency components.